OpenStudy (anonymous):

X school purchased 357 pen and 323 rubber and 289 pencils.The total bill amounts to 4054.5.On the other day school Y purchased 221 pens,247 rubbers and 273 pencils which amounts to 3074.5.Find the price of 100 pens ,100 rubbers and 100 pencils

OpenStudy (anonymous):

Let the pen be x , rubber be y and pencils be z: 357x + 323y + 289z = 4054.5 221x + 247y + 273z = 3074.5 How can you find three variables from two equations??

OpenStudy (anonymous):

Let the cost of 1 pen , 1 pencil, 1 rubber be x, y,z 357x+ 289y+ 323z =4054.5

OpenStudy (anonymous):

only possibility for solving is add or subtracting 2 equations but it does'nt work:(

OpenStudy (anonymous):

Yes.. It is unsolvable in my opinion..

OpenStudy (anonymous):

its a question of one of the competative exams CAT